OpenStudy (samanthagreer):

Combine like terms. 8 + 6x – 11x + 12 A. 20 + 5x B. 20 – 17x C. 14x + 1x D. 20 – 5x

OpenStudy (anonymous):

8+12=20 6x-11x=-5x

OpenStudy (anonymous):

Combine like terms for 8 + 6x–11x + 12 Evaluate the x terms: 6x + 11x (6 + 11)x 17x Evaluate the constants: 8 + 12 20 Combining all like terms, we get: 17x + 20 -

OpenStudy (catlover5925):

20-5x or D

OpenStudy (samanthagreer):

thanks guys

OpenStudy (samanthagreer):

can u help me with another

OpenStudy (anonymous):

@samanthagreer it is B because Combine like terms for 8 + 6x–11x + 12 Evaluate the x terms: 6x + 11x (6 + 11)x 17x Evaluate the constants: 8 + 12 20 Combining all like terms, we get: 17x + 20 - And you can flip them because it is adding, here is the link ; http://www.mathcelebrity.com/liketerm.php?num=8+%2B+6x+%E2%80%93+11x+%2B+12&pl=Combine+Like+Terms

OpenStudy (samanthagreer):

A. –10 B. –14 C. –3 D. 1

OpenStudy (anonymous):

It has to be -14 because -4 times 3 is -12 and it says + 2 so -12 + 2 + 14

OpenStudy (samanthagreer):

Evaluate the expression by using the given value of the variables. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16 A. 21 B. 36 C. –3 D. 3

OpenStudy (anonymous):

y/2 (4)− 10/x +6 16/2 (4)−10/5 +6 =36